• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ramesh/Vince

    Log entry time 03:19:22 on November 05, 2009

    Entry number 297453

    keyword=controlled access work in Left HRS

    Vince and I took controlled access to troubleshoot the trigger issue with
    the left HRS (Diancheng and Xiaoyan went to the right HRS for other work
    in the same time). We found that the retiming signal was before the L1A
    (which is against the prescription that retiming needs to be about 40 to
    60 ns later than L1A). Diancheng recalls that the prescription was
    followed and retiming was set that way; not sure why we see retiming
    before the L1A now. We set the retiming 60 ns delayed with respect to
    L1A. Took a new run and found all histograms normal. We were even more
    happier that the double peak structure in S1 and S2 TDCs now disappears;
    no second T0 in VDC too. Everything looking normal, we called off
    controlled access to go to beam permit.